Tim can fold a certain number of brochures in 20 minutes. John can fold the same number in 15 minutes. If tim and john work together for 3 hours and during that time john folds x brochures, how many brochures will tim fold?
A)4/3x
B)5/4x
C)3/4x
D)2/3x
E)1/2x

Answer by richwmiller(17219)

1/4vs 1/3
1/4/1/3=4/3
A) 4/3 x
or
john 3/1/4=12
tim 3/1/3=9
j/t=12/9=4/3
A)4/3 x